solve the equation [0 0 -1]
[3 x 0]
[2 0 3] = x 2

Answer by stanbon(75887)

solve the equation [0 0 -1]
[3 x 0]
[2 0 3] = x 2
------------------
[0+0+0-[-2x+0+0]] = x^2
---
2x = x^2
---
x^2-2x = 0
x(x-2) = 0
x = 0 or x = 2
=====================
Cheers,
Stan H.
============